# FGBERT: Function-Driven Pre-trained Gene Language Model for Metagenomics

### v1

@@ Appendix G Detailed Classification Results on CARD Dataset
Table 9: Detailed Classification Results of FGBERT Model on AMR Gene Family Category Prediction. Our model performs a three-class classification prediction task for each gene sequence in the CARD dataset, covering AMR Gene Family (269 classes in total), Drug Class (37 classes in total) and Resistance Mechanism (7 classes in total). In order to comprehensively evaluate the model performance, we used several evaluation metrics, including Accuracy, Macro Pre, Macro Recall, Macro F1, Weighted Pre, Weighted Recall and Weighted F1. One of the challenging tasks is to predict the AMR Gene Family category, which has up to 269 different classes and requires high complexity classification of gene sequences with multiple labels. This requires the model to be able to understand the biological features of the gene sequences and accurately map them to the corresponding AMR Gene Family categories to ensure accurate prediction in the multi-classification task.

### v2

@@ D.2 Detailed Classification Results on CARD Dataset
Table 19: Detailed Classification Results of FGBERT on AMR Gene Family Category Prediction.
@@
Appendix E Broader Impacts
In our work, we focus on the wider application of macro-genomic pre-training models, particularly in the field of functional annotation. We argue that the positive societal impacts of such models include, but are not limited to, improved disease surveillance and prediction, facilitation of new drug development and optimisation of vaccine design. However, we should also be aware of possible negative impacts such as potential risk of misuse and privacy concerns.
Appendix F Prelinaries.
$\textit{Attention}(Q,K)=\frac{Q\cdot K^{T}}{\sqrt{d_{k}}}$ (7)
$Z=\textit{softmax}(\textit{Attention}(Q,K))V$ (8)